# Time multidimensional Markov Renewal chains - An algebraic approach

### v1

@@ 2.1 Preliminaries
$[A*B]\left(k_{1:d}\right):=\sum_{l+l^{\prime}=k_{1:d}}A(l)B(l^{\prime}),\quad k_{1:d}\in\mathbb{N}^{d}.$ (2.1)
@@
In the above definition the emphasis is given on the representation of $A*B$ as a multidimensional sequence of matrices. Nevertheless, it is also useful to interpret $A*B$ as a matrix of multidimensional sequences. Indeed, notice by (2.1) that
$[A*B]_{ij}\left(k_{1:d}\right)=\sum_{r\in E}\sum_{l+l^{\prime}=k_{1:d}}\alpha_{ir}(l)\beta_{rj}(l^{\prime}),\quad i,j\in E,\,k_{1:d}\in\mathbb{N}^{d},$
thus $A*B=\left([A*B]_{ij}\right)_{i,j\in E}$, where $[A*B]_{ij}=\sum_{r\in E}\alpha_{ir}*\beta_{rj}$.
@@
The computational complexity of the convolution product is an important practical issue to be addressed. Using the exact formula to compute $[A*B](l_{1:d})$ for all $d$-tuples of indices $0\leq l_{1:d}\leq k_{1:d}$ has $\mathcal{O}(s^{3}\,k^{2}_{1}\cdots k^{2}_{d})$ complexity. An effective way to reduce this complexity is to use the Fast Fourier Transform (FFT)[[7]] for matrix sequences.
@@
$[\mathrm{dg}(\mathbbm{1}_{s})*A](k_{1:d})=\sum_{l=0_{d}}^{k_{1:d}}A(l),\quad k_{1:d}\in\mathbb{N}^{d}.$
@@ 2.2 A representation of the convolutional inverse
Indeed, if the latter holds, then the expression of $A^{(-1)}$ in the right side of the first equality of (2.3) follows from the fact that $A(k_{1:d})=A_{0}(k_{1:d})A(0_{d})$, where $A_{0}(k_{1:d})$ is given by (2.4), and $A_{0}$ satisfies (2.6). Let us now proceed to the proof of (2.6). Since $[\mathbbm{I}_{s}-A](0_{d})=O_{s}$, by Remark (2.3) the series below is well defined and
@@ 2.3 Inversion with the classical recurrence
This formulation enables the computation of each component of $A^{(-1)}$, iteratively, using previously computed terms and an additional multiplication with the inverse of $A(0_{d})$. The resulting complexity of the above method is $\mathcal{O}\left(s^{3}\prod_{i=1}^{d}k^{2}_{i}\right)$.
@@ 2.4 Newton’s inversion method
i.e., $b(x_{1:d})$ is the multiplicative inverse of $a(x_{1:d})$ or using the isomorphism in Theorem 2.1, the coefficients $g$ form the convolutional inverse of $f$. The performance of this computation can be significantly improved using Newton’s iteration to achieve high precision very quickly. First, we truncate (2.8) to a finite degree and take the approximation $b_{k_{1:d}}$ of $b$ such that
$a\cdot b_{k_{1:d}}=1\mod\left<x_{1}^{2{k_{1}}},\ldots,x_{d}^{2^{k_{d}}}\right>.$
@@
$b_{k_{1:d}+1_{s}}=b_{k_{1:d}}+b_{k_{1:d}}\cdot(2-a\cdot b_{k_{1:d}})\mod\left<x_{1}^{2{k_{1}+1}},\ldots,x_{d}^{2^{k_{d}+1}}\right>,$ (2.9)
where $1_{s}$ refers to an $s$-dimensional vector of ones. The proof of (2.9) is deferred to the Appendix B. The corresponding complexity, when using FFT, is reduced to $\mathcal{O}\left(\prod_{i=1}^{d}k_{i}\log_{2}(k_{i})\right).$
@@
$B_{k_{1:d}+1_{s}}=B_{k_{1:d}}(2\cdot I_{s}-A\cdot B_{k_{1:d}})\mod\left<x^{2^{k_{1}+1}}_{1},\ldots,x^{2^{k_{d}+1}}_{d}\right>,$ (2.10)
with initial condition $B_{0_{d}}=[A(0_{d})]^{-1}$. This formulation allows us to double the degree of approximation at each iteration, making the method particularly suitable for large-scale computations. As a result, the computational complexity for calculating the convolutional inverse is reduced to $\mathcal{O}\left(s^{3}\prod_{i=1}^{d}k_{i}\log_{2}(k_{i})\right).$
@@ 2.5 Inversion with the Gauss-Jordan algorithm
Let $A\in\mathcal{M}_{s}(\mathbb{N}^{d})$. Performing any elementary row operation on $A$ corresponds to the convolution $E*A$, where $E$ is an appropriate elementary matrix sequence derived by $\mathbbm{I}_{s}$.
@@
    $A$ is row equivalent to the identity sequence $\mathbb{I}_{s}$.
@@
    $A$ can be written as the convolution of a finite sequence of elementary matrix sequences.
@@
[…]
@@ 3 Multi-time Markov renewal chains
[…]
@@
[…]
@@
[…]
@@
Let $(J,S)$ be a $dt$–MRC. For any $i\in E$, $k_{1:d}\in\mathbb{N}^{d}$ and $k_{u}\in\mathbb{N},\ u\in[1;d]$, we introduce the following functions related to the sojourn time distribution of $[X_{n+1}|J_{n}=i],\ n\in\mathbb{N}$:
@@
[…]
@@
Using the Markov renewal property of the sequence $(J,\,S)$, we conclude that the events
@@
If $(J,\,S)$ is a dt-MRC, then the process $(J,\,S^{[y]})$ with interjump times $Y_{n+1}=X^{[u]}_{n+1}\cdot X^{[v]}_{n+1}$,  $1\leq u<v\leq d$, forms a Markov renewal chain and its semi-Markov kernel $q^{[y]}$ can be determined by:
@@
For the MRC $(J,S^{[y]})$ the marginal survival function $\overline{H}^{[y]}_{i}(\cdot)$ of state i, is given by
@@
Since $q(0_{d})$ is assumed to be the null matrix then $[\mathbbm{I}_{s}-q](0_{d})=I_{s}$ and consequently the convolutional inverse of the sequence of matrices $\mathbbm{I}_{s}-q$ exists by Theorem 2.1. Furthermore, by (2.3) we get directly that
@@
[…]
@@
[…]
@@
[…]
@@ 3. $(iii)$
[…]
@@ 4.1 Performing Convolution
For the FFT-based method, we processed two matrix sequences up to time (128,128)(128,128)( 128 , 128 ), whereas for the direct method, the computation was limited to (32,32)(32,32)( 32 , 32 ) due to its higher computational burden.
@@
Gauss-Jordan- up to (128,128)(128,128)( 128 , 128 ) 100 38.81 1.000 35.96 2.80
Direct-up to (32,32)(32,32)( 32 , 32 ) 100 80.56 2.076 80.30 0.14
@@ 4.2 Moments of MRC
[…]
@@
    $f_{i\bullet}(x_{1:2})=\exp{\left\{-\left(\alpha_{i}+\beta_{i}+\gamma_{i}\right)\right\}}\frac{\alpha^{x_{1}}_{i}}{(x_{1})!}\frac{\beta^{x_{2}}_{i}}{(x_{2})!}\sum_{k=0}^{\min(x_{1},x_{2})}\binom{x_{1}}{k}\binom{x_{2}}{k}k!\left(\frac{\gamma_{1}}{\alpha_{i}\cdot\beta_{i}}\right)$
  • First moments of the first sojourn time dimension: $m^{[1]}_{i}$=$\alpha_{i}+\gamma_{i}+1$.
  • First moments of the second sojourn time dimension: $m^{[2]}_{i}$=$\beta_{i}+\gamma_{i}+1$.
@@
[…]
@@ Appendix A Fast Fourier Transform
The discrete time Fourier transform (DFT) of a sequence $A$ $\in\mathcal{M}_{s}(\mathbb{N}^{d})$ is a complex matrix valued function defined by
@@
Let $\widehat{A}$ be the DFT of a sequence $A$ $\in\mathcal{M}_{s}(\mathbb{N}^{d})$. The inverse DFT is determined by
@@
    Take the convolution $A*B$ as the final output.
@@ Appendix B Newton’s inversion method proof
$e_{k_{1:d}}:=1-a\cdot b_{k_{1:d}}\in\left<x_{1}^{2^{k_{1}}},\ldots,x_{d}^{2^{k_{d}}}\right>,$
since $a\cdot b_{k_{1:d}}=1\mod\left<x_{1}^{2^{k_{1}}},\ldots,x_{d}^{2^{k_{d}}}\right>$. Now, if $y:=b_{k_{1:d}}+b_{k_{1:d}}\cdot\alpha$, then we have
$y\cdot a$ $=$ $\left(b_{k_{1:d}}+b_{k_{1:d}}\cdot e_{k_{1:d}}\right)\cdot a=b_{k_{1:d}}\cdot a+b_{k_{1:d}}\cdot e_{k_{1:d}}\cdot a$
$=$ $1-e_{k_{1:d}}+b_{k_{1:d}}\cdot e_{k_{1:d}}\cdot a=1-e_{k_{1:d}}\cdot(1-b_{k_{1:d}}\cdot a)$
$=$ $1-e_{k_{1:d}}^{2}.$
Since $e_{k_{1:d}}\in\left<x_{1}^{2^{k_{1}}},\ldots,x_{d}^{2^{k_{d}}}\right>$, it follows that $e^{2}_{k_{1:d}}\in\left<x_{1}^{2^{k_{1}+1}},\ldots,x_{d}^{2^{k_{d}+1}}\right>.$ Hence,
$y\cdot a=1\mod\left<x_{1}^{2^{k_{1}+1}},\ldots,x_{d}^{2^{k_{d}+1}}\right>,$
as required. From the uniqueness of the power series inverse, we conclude that $y$ contains the elements of the convolutional inverse up to $(2^{k_{1}+1},\ldots,2^{k_{d}+1})$.

### v4

@@ 2.1 Preliminaries
$A*B]\left(k_{1:d}\right):=\sum_{l+l^{\prime}=k_{1:d}}A(l)B(l^{\prime}),\quad k_{1:d}\in\mathbb{N}^{d}.$ (2.1)
@@
In the above definition the emphasis is given on the representation of ABA*B as a multidimensional sequence of matrices. Nevertheless, it is also useful to interpret ABA*B as a matrix of multidimensional sequences. Indeed, notice by (2.1) that
$A*B]_{ij}\left(k_{1:d}\right)=\sum_{r\in E}\sum_{l+l^{\prime}=k_{1:d}}\alpha_{ir}(l)\beta_{rj}(l^{\prime}),\quad i,j\in E,\,k_{1:d}\in\mathbb{N}^{d},$
thus $A*B=\left([A*B]_{ij}\right)_{i,j\in E}$, where $A*B]_{ij}=\sum_{r\in E}\alpha_{ir}*\beta_{rj}$.
@@
The computational complexity of the convolution product is an important practical issue to be addressed. Using the exact formula to compute $A*B](l_{1:d})$ for all $d$-tuples of indices $0\leq l_{1:d}\leq k_{1:d}$ has $\mathcal{O}(s^{3}\,k^{2}_{1}\cdots k^{2}_{d})$ complexity. An effective way to reduce this complexity is to use the Fast Fourier Transform (FFT)[[7]] for matrix sequences.
@@
$\mathrm{dg}(\mathbbm{1}_{s})*A](k_{1:d})=\sum_{l=0_{d}}^{k_{1:d}}A(l),\quad k_{1:d}\in\mathbb{N}^{d}.$
@@ 2.2 A representation of the convolutional inverse
Indeed, if the latter holds, then the expression of $A^{(-1)}$ in the right side of the first equality of (2.3) follows from the fact that $A(k_{1:d})=A_{0}(k_{1:d})A(0_{d})$, where $A_{0}(k_{1:d})$ is given by (2.4), and $A_{0}$ satisfies (2.6). Let us now proceed to the proof of (2.6). Since $\mathbbm{I}_{s}-A](0_{d})=O_{s}$, by Remark (2.3) the series below is well defined and
@@ 2.3 Inversion with the classical recurrence
This formulation enables the computation of each component of $A^{(-1)}$, iteratively, using previously computed terms and an additional multiplication with the inverse of $A(0_{d})$. The resulting complexity of the above method is $\mathcal{O}\left(s^{3}\prod_{i=1}^{d}k^{2}_{i}\right)$. Therefore, if we need to compute the convolutional inverse up to time $(2^{N},\ldots,2^{N})$, the complexity becomes to $\mathcal{O}\left(s^{3}\cdot 4^{N\cdot d}\right)$.
@@ 2.4 Newton’s inversion method
i.e., $b(x_{1:d})$ is the multiplicative inverse of $a(x_{1:d})$ or using the isomorphism in Theorem 2.1, the coefficients $g$ form the convolutional inverse of $f$. The performance of this computation can be significantly improved using Newton’s iteration to achieve high precision very quickly. First, we truncate (2.8) to a finite degree and take the approximation $b_{N}$ of $b$ such that
$a\cdot b_{N}=1\mod\left<x_{1},\ldots,x_{d}\right>^{2^{N}}.$
@@
$b_{N+1}=b_{N}+b_{N}\cdot(2-a\cdot b_{N})\mod\left<x_{1},\ldots,x_{d}\right>^{2^{N+1}},$ (2.9)
The proof of (2.9) is deferred to the Appendix B. The corresponding complexity up to time $(2^{N},\ldots,2^{N})$, when using FFT, is reduced to $\mathcal{O}\left(N^{d}\cdot 2^{d\cdot N}\right)$.
@@
$B_{N+1}=B_{N}+B_{N}\cdot(2-A\cdot B_{N})\mod\left<x_{1},\ldots,x_{d}\right>^{2^{N+1}},$ (2.10)
with initial condition $B_{0_{d}}=[A(0_{d})]^{-1}$. This formulation allows us to double the degree of approximation at each iteration, making the method particularly suitable for large-scale computations. As a result, the computational complexity for calculating the convolutional inverse is reduced to $\mathcal{O}\left(s^{3}\prod_{i=1}^{d}k_{i}\log_{2}(k_{i})\right).$ As a result, the overall computational complexity of this procedure up to time $(2^{N},\ldots,2^{N})$ is $\mathcal{O}\left(s^{3}\cdot N^{d}\cdot 2^{d\cdot N}\right)$.
@@ 2.5 Inversion with the Gauss-Jordan algorithm
Let $A\in\mathcal{M}_{s}(\mathbb{N}^{d})$. Performing any elementary row operation on $A$ corresponds to the convolution EAE*A, where $E$ is an appropriate elementary matrix sequence derived by $\mathbbm{I}_{s}$.
@@
    AA is row equivalent to the identity sequence $\mathbb{I}_{s}$.
@@
    AA can be written as the convolution of a finite sequence of elementary matrix sequences.
@@
$\begin{pmatrix}\mathbb{I}_{1}\ &\ b*\alpha^{(-1)}\\ 0\ &\ \mathbb{I}_{1}\end{pmatrix}$ $\xrightarrow{R_{21}[-b*\alpha^{(-1)}]}{}$ $\begin{pmatrix}\mathbb{I}_{1}\ &\ 0\\ 0\ &\ \mathbb{I}_{1}\end{pmatrix}$.
@@ 3 Multi-time Markov renewal chains
then one gets directly that for $1\leq u\leq d$, the matrix-valued function $q^{[u]}$ is a $1t$-smk, possibly allowing instantaneous transitions. The same holds true for any other simultaneous selection of $d_{0}<d$ time components, thus forming a $d_{0}t$-smk.
@@
Let $(J,S)$ be a dt-MRC and $q$ its associated dt-smk. Then, the processes $(J,S)$, $(J,X)$, $(J,S^{[u]})$ and $(J,X^{[u]})$, $u\in[1;d]$, are Markov chains with transition probabilities:
@@
where $q^{[u]}_{ij}(k_{u})$ is defined in (3.2). The chain $(J,S^{[u]})$ is a 1t-MRC with associated smk $q^{[u]}$. The process $J$ is also a Markov chain, called the embedded Markov chain associated to the dt-MRC (J,S)(J,\,S), as well as to the 1t-MRC $(J,S^{[u]})$. Its transition probabilities are given by
@@
Let $(J,S)$ be a dtdt–MRC. For any $i\in E$, $k_{1:d}\in\mathbb{N}^{d}$ and $k_{u}\in\mathbb{N},\ u\in[1;d]$, we introduce the following functions related to the sojourn time distribution of $X_{n+1}|J_{n}=i],\ n\in\mathbb{N}$:
@@
Let $(J,S)$ be a dtdt–MRC and $(J,S^{[u]})$ the associated marginal 1t1t–MRCs, $u\in[1;d]$. For any $i,j\in E$, $k_{1:d}\in\mathbb{N}^{d}$ and $k_{u}\in\mathbb{N},\ u\in[1;d]$, we introduce the following functions related to the conditional sojourn time distribution of $X_{n+1}|J_{n}=i,J_{n+1}=j],\ n\in\mathbb{N}$:
@@
Using the Markov renewal property of the sequence (J,S)(J,\,S), we conclude that the events
@@
If (J,S)(J,\,S) is a dt-MRC, then the process $(J,\,S^{[y]})$ with interjump times $Y_{n+1}=X^{[u]}_{n+1}\cdot X^{[v]}_{n+1}$,  $1\leq u<v\leq d$, forms a Markov renewal chain and its semi-Markov kernel $q^{[y]}$ can be determined by:
@@
[…]
@@
Since $q(0_{d})$ is assumed to be the null matrix then $\mathbbm{I}_{s}-q](0_{d})=I_{s}$ and consequently the convolutional inverse of the sequence of matrices $\mathbbm{I}_{s}-q$ exists by Theorem 2.1. Furthermore, by (2.3) we get directly that
@@
is said to be the $d$-dimensional multi-time semi-Markov chain (dt-SMC) associated to $(J,S)$.
@@
is called the time $d$-dimensional Markov renewal function of the dt-MRC $(J,S)$.
@@
Let $(J,S)$ be a mt-MRC and $(J,S^{[u]})$, $1\leq u\leq d$, the associated marginal Markov renewal chains. The process $(J,S)$ is said to be
@@ 3. $(iii)$
    $r_{1:d}$–periodic if the chain $(J,S^{[u]})$ is periodic with period $r_{u}$, for $1\leq u\leq d$. If additionally, $r_{u}=r$, for $1\leq u\leq d$, then $(J,S)$ is said to be simply rr–periodic and an 11–periodic mt-MRC is also referred to as aperiodic,
@@ 4.1 Performing Convolution
For the FFT-based method, we processed two matrix sequences up to time $(128,128)$, whereas for the direct method, the computation was limited to $(32,32)$ due to its higher computational burden.
@@
Gauss-Jordan- up to $(128,128)$ 100 38.81 1.000 35.96 2.80
Direct-up to $(32,32)$ 100 80.56 2.076 80.30 0.14
@@ 4.2 Moments of MRC
Let us consider a three-state 2-t Markov renewal chain $(J,S)$, where the embedded Markov chain $J$ with state space $\{1,2,3\}$, has the following transition matrix
@@
    $f_{i\bullet}(x_{1:2})=\exp{\left\{-\left(\alpha_{i}+\beta_{i}+\gamma_{i}\right)\right\}}\frac{\alpha^{x_{1}}_{i}}{(x_{1})!}\frac{\beta^{x_{2}}_{i}}{(x_{2})!}\sum_{k=0}^{\min(x_{1},x_{2})}\binom{x_{1}}{k}\binom{x_{2}}{k}k!\left(\frac{\gamma_{i}}{\alpha_{i}\cdot\beta_{i}}\right)^{k}$
  • First moments of the first sojourn time dimension: $m^{[1]}_{i}$$\alpha_{i}+\gamma_{i}+1$.
  • First moments of the second sojourn time dimension: $m^{[2]}_{i}$$\beta_{i}+\gamma_{i}+1$.
@@
$q=\left({\begin{array}[]{ccc}0&\ \frac{1}{3}\cdot f_{12}&\ \ \frac{2}{3}\cdot f_{13}\\ &&\\ \frac{1}{5}\cdot f_{21}&\ 0&\ \ \frac{4}{5}\cdot f_{23}\\ &&\\ \frac{3}{4}\cdot f_{31}&\ \ \frac{1}{4}\cdot f_{32}&\ \ 0\end{array}}\right),$
@@ Appendix A Fast Fourier Transform
The discrete time Fourier transform (DFT) of a sequence AA $\in\mathcal{M}_{s}(\mathbb{N}^{d})$ is a complex matrix valued function defined by
@@
Let $\widehat{A}$ be the DFT of a sequence AA $\in\mathcal{M}_{s}(\mathbb{N}^{d})$. The inverse DFT is determined by
$A_{ij}(k_{1:d})=\frac{1}{(2\pi)^{d}}\int_{[-\pi,\,\pi]^{d}}\widehat{A}_{ij}\left(e^{\boldsymbol{i}\,\omega_{1:d}}\right)\exp\{\boldsymbol{i}\left(k_{1:d}\cdot\omega_{1:d}\right)\}d\omega_{1:d},\quad k_{1:d}\in\mathbb{N}^{d},\,i,j\in E.$
@@
    Take the convolution ABA*B as the final output.
@@ Appendix B Newton’s inversion method proof
$e_{N}:=1-a\cdot b_{N}\in\left<x_{1},\ldots,x_{d}\right>^{2^{N}},$
since $a\cdot b_{N}=1\mod\left<x_{1},\ldots,x_{d}\right>^{2^{N}}$. Now, if $y:=b_{N}+b_{N}\cdot\alpha$, then we have
$y\cdot a$ $=$ $\left(b_{N}+b_{N}\cdot e_{N}\right)\cdot a=b_{N}\cdot a+b_{N}\cdot e_{N}\cdot a$
$=$ $1-e_{N}+b_{N}\cdot e_{N}\cdot a=1-e_{N}\cdot(1-b_{N}\cdot a)$
$=$ $1-e_{N}^{2}.$
Since $e_{N}\in\left<x_{1},\ldots,x_{d}\right>^{2^{N}}$, it follows that $e^{2}_{N}\in\left<x_{1},\ldots,x_{d}\right>^{2^{N+1}}.$ Hence,
$y\cdot a=1\mod\left<x_{1},\ldots,x_{d}\right>^{2^{N+1}},$
as required. From the uniqueness of the power series inverse, we conclude that $y$ contains the elements of the convolutional inverse up to $(2^{N+1},\ldots,2^{N+1})$.